# DFDT: Dynamic Fast Decision Tree for IoT Data Stream Mining on Edge Devices

### v1

@@ 5 Conclusions
In this paper, the Dynamic Fast Decision Tree (DFDT) was introduced for memory-constrained data stream mining. DFDT consolidates different adaptive strategies from prior research to control decision tree growth effectively. These features make DFDT particularly suitable for resource-constrained environments, such as IoT devices, edge computing applications, and real-time monitoring systems. Experiments show the proposed consolidated adaptive strategies achieve improved predictive performance (0.43 vs. 0.29 ranking) compared to VFDT and SVFDT, with significantly reduced runtime under 12 diverse datasets.
Despite its strengths, DFDT has limitations, particularly in handling non-stationary data streams. While it inherits VFDT’s robust response to concept drifts through resplitting nodes and approximating information gains [14], it could be further improved by incorporating adaptive strategies, such as memory-conservative alternate subtree growth when a concept drift is detected at a node, using loss-based detection mechanisms like ADWIN [2]. Addressing such limitations could enhance DFDT’s adaptability to diverse, rapidly changing environments. Additionally, testing DFDT on a broader range of datasets, including noisy imbalanced high-dimensional data streams, could provide further insights into its robustness.
@@ Acknowledgments
This work was funded by the EU, through the Portuguese Republic’s Recovery and Resilience Plan, within the project PRODUTECH R3. It was also funded by the Portuguese Foundation for Science and Technology under project doi.org/10.54499/UIDP/00760/2020 and Ph.D. scholarship PRT/BD/154713/2023.

### v2

@@ Conclusions
Dynamic Fast Decision Tree (DFDT) combines different adaptive strategies from prior research to control decision tree growth. These features make DFDT particularly suitable for resource-constrained environments, such as IoT devices, edge computing applications, and real-time systems.
Extensive experimental evaluation, including ablation studies and statistical comparisons across a diverse set of data streams, demonstrated that DFDT variants consistently outperform or match baseline learners. In particular, $\text{DFDT}_{\text{Low}}$ was identified as the most resource-efficient configuration, achieving strong performance with minimal resource consumption. $\text{DFDT}_{\text{Medium}}$ offered a favorable trade-off between predictive accuracy and efficiency, while $\text{DFDT}_{\text{High}}$ achieved the highest accuracy overall, justifying increased memory and runtime costs when accuracy is the primary concern.
Unlike traditional VFDT-based learners, DFDT incorporates fine-grained control over growth and memory usage without sacrificing model adaptability. Moreover, its compatibility with existing ensemble frameworks allows it to serve as a drop-in replacement for VFDT, offering immediate benefits in large-scale, real-time stream mining applications. Future work may explore the integration of DFDT in different ensemble frameworks and investigate its performance on a broader range of datasets, including noisy imbalanced high-dimensional data streams.
@@ Acknowledgments
Work funded by Portuguese Foundation for Science and Technology under Ph.D. scholarship PRT/BD/154713/2023 and project doi.org/10.54499/UIDP/00760/2020.